OpenStudy (anonymous):

What is the height of a square based prism whose base edge is 6cms and lateral surface area is 288sq.cms

OpenStudy (whpalmer4):

 Lateral surface area = 4*area of one side = 4*l*w = 4*6*h \[4*6*h = 288\]\[h=288/24 = ?\]

OpenStudy (anonymous):

12

OpenStudy (anonymous):

that is the answer

OpenStudy (anonymous):

thank u @whpalmer4

OpenStudy (whpalmer4):

You're welcome!